# Constacyclic codes with best-known parameters

### v1

@@ 6 Constacyclic codes with $n=\Phi_{p^{b}}(q)$
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-p}{p^{b}}.$
Moreover, noticing $pN_{p}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{p,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorems 6.1, 6.2 and 6.3, if furthermore $r\leq 2$, then by Theorem 3.3, we can get a better lower bound on minimum distance for the second family of constacyclic codes.
In Table LABEL:tab:pb+1c, we present some codes constructed in Section 6 having good parameters.
@@
$q$ $p$ $b$ $r$ Lower Bose Parameter Optimality
bound distance
22 22 33 11 22 33 $17,8,6]^{*}$ optimal
22 22 33 11 22 33 $[17,9,5]$ optimal
22 33 22 11 88 $12$ $73,36,14]^{*}$ $d_{\text{best}}=16$
22 33 22 11 88 $11$ $[73,37,13]$ $d_{\text{best}}=14$
33 22 22 11 22 33 $10,5,4]^{*}$ optimal
33 22 22 22 22 66 $10,4,6]^{*}$ optimal
33 22 22 22 11 33 $[10,6,4]$ optimal
33 22 33 22 88 99 $82,40,18]^{*}$ $d_{\text{best}}=21$
44 22 22 11 33 44 $17,8,6]^{*}$ $d_{\text{best}}=8$
44 22 22 33 33 77 $17,8,8]^{*}$ best-known
44 22 22 33 22 77 $[17,9,7]$ best-known
55 22 22 22 44 55 $26,13,8]^{*}$ $d_{\text{best}}=10$
55 22 22 44 44 99 $26,12,10]^{*}$ $d_{\text{best}}=11$
55 22 22 44 33 44 $[26,14,8]$ $d_{\text{best}}=9$
77 22 22 11 77 $10$ $50,25,18]^{*}$ best-known
77 22 22 22 77 99 $[50,26,14]$ $d_{\text{best}}=15$
88 22 22 11 $10$ $22$ $65,32,22]^{*}$ best-known
@@ 7 Constacyclic codes with $n=\Phi_{p_{1}p_{2}}(q)$
Recall that $q$ is a prime power, and $r\mid q-1$. Besides, we assume that $p_{1},p_{2}$ are two distinct primes with $p_{1}<p_{2}$. In the following, we construct some infinite families of constacyclic codes with
@@
Moreover, we only consider the case when the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$, has exactly two elements, and then we use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
@@
Let $p_{1},p_{2}$ be two distinct primes with $p_{1}<p_{2}$, and $n=\frac{q^{p_{1}p_{2}}-1}{\mathrm{lcm}(q^{p_{1}}-1,q^{p_{2}}-1)}$. Assume $p_{2}\nmid(q^{p_{1}}-1)/r$. Then, an infinite family of constacyclic $n,n-1-p_{1}p_{2}\lceil(n-1)/(2p_{1}p_{2})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p_{1}p_{2})\rceil}{q-1}\right\rceil+\Lambda_{p_{1}p_{2}}]$ codes is constructed, where $\Lambda_{p_{1}p_{2}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p_{1}p_{2}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p_{1}p_{2}\lfloor(n-1)/(2p_{1}p_{2})\rfloor,\geq\lfloor\frac{q(n-1)}{2p_{1}p_{2}(q-1)}\rfloor]$ codes is constructed.
@@
First, we want to determine $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. Since $r\mid q-1$, we write
$nr=\frac{q^{p_{1}p_{2}}-1}{\mathrm{lcm}(q^{p_{1}}-1,q^{p_{2}}-1)/r},$ (7.1)
where the denomenator is a positive integer. Combining (7.1) with Lemma 2.1, we have
$l_{i}\mid p_{1}p_{2}\quad\textrm{for}\quad i\in Z_{n,r}.$
@@
$\equiv 1+1+\cdots+1+1\equiv p_{2}\pmod{(q^{p_{1}}-1)/r}$.
@@
$\gcd\left(\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1},\frac{q^{p_{1}}-1}{r}\right)=\gcd\left(p_{2},\frac{q^{p_{1}}-1}{r}\right).$
So, noticing $\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1}=n\cdot\frac{q^{p_{2}}-1}{q-1}$ and $\gcd(p_{2},q^{p_{2}}-1)=1$, we obtain
$\gcd\left(n,\frac{q^{p_{1}}-1}{r}\right)=\gcd\left(p_{2},\frac{q^{p_{1}}-1}{r}\right).$ (7.4)
Combining this with the assumption $p_{2}\nmid(q^{p_{1}}-1)/r$, we have
$\gcd\left(n,\frac{q^{p_{1}}-1}{r}\right)=1.$ (7.5)
@@
Next, we want to compute $N_{p_{2}}$. Similarly as (7.4), we have
$\gcd\left(n,\frac{q^{p_{2}}-1}{r}\right)=\gcd\left(p_{1},\frac{q^{p_{2}}-1}{r}\right).$
@@
Hence, we have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p_{1}p_{2}\}.$
Then, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that the condition $p_{2}\nmid(q^{p_{1}}-1)/r$ in Theorem 7.1 holds when $p_{1}\nmid p_{2}-1$. Indeed, if $p_{1}\nmid p_{2}-1$, then $q^{p_{1}}\not\equiv 1\pmod{p_{2}}$, and so $p_{2}\nmid q^{p_{1}}-1$.
We also remark that from (7.4) and the computation of $N_{p_{1}}$, one can see that $N_{p_{1}}=0$ if and only if $p_{2}\nmid(q^{p_{1}}-1)/r$. This means that the set $\{l_{i}:\,i\in Z_{n,r}\}$ has exactly two elements if and only if $p_{2}\nmid(q^{p_{1}}-1)/r$.
In addition, in Theorem 5.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{q(n-1)}{2p_{1}p_{2}(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor(n-1)/(2p_{1}p_{2})\rfloor}{q-1}\right\rceil+\Lambda_{p_{1}p_{2}}^{\prime},$
where $\Lambda_{p_{1}p_{2}}^{\prime}$ has been defined in Theorem 3.3.
In Table LABEL:tab:thmp1p2f, we present some codes constructed in Theorem 7.1 having good parameters.
@@
$q$ $p_{1}$ $p_{2}$ $r$ Lower Bose Parameter Optimality
bound distance
@@
77 22 33 66 44 66 $[43,25,11]$ $d_{\text{best}}=12$

### v2

@@ 5 Constacyclic codes with $n=\Phi_{p^{b}}(q)$
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}=\frac{n-p}{p^{b}}.$
Moreover, noticing $pN_{p}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have $\{l_{i}:\,i\in Z_{n,r}\}=\{p,p^{b}\}.$ Then, the desired result follows directly from Theorems 3.1 and 3.4. ∎
We remark that in Theorem 5.3, from (5.7) and (5.9) we have $\gcd(n,q-1)=p$, which does not divide (q1)/r(q-1)/r by assumption; and so, the sufficient condition in Lemma 2.18 about equivalence between constacyclic codes and cyclic codes fails in this case.
We notice that some codes in this section are better than BCH constacyclic codes; see Example 5.4.
Example 5.4.
Let $q=4$, $p=2$, $b=2$, $r=3$, and $n=17$. Then, the $4$-cyclotomic cosets in $Z_{17,3}$ are: $C_{1}=\{1,4,13,16\}$, $C_{7}=\{7,10,28,40\}$, $C_{19}=\{19,25,43,49\}$, $C_{22}=\{22,31,37,46\}$, $C_{34}=\{34\}$. So, $\{l_{i},i\in Z_{17,3}\}=\{1,4\}$, $N_{1}=1$ and $N_{4}=4$. From the first family of constacyclic codes in Theorem 5.1, we take the defining set as the set $C_{1}\cup C_{7}\cup C_{34}$ and then obtain a $4$-ary constacyclic $[17,8,8]$ code, which is not a BCH code and has the best-known parameters [18]. In this case, there is only one BCH constacyclic code of length 17 and dimension 8 (its defining set is $C_{7}\cup C_{22}\cup C_{34}$ and minimum distance is 6).
In Table LABEL:tab:qpb, we present some codes constructed in Section 5 having good parameters. In the table, $d_{\text{best}}$ stands for the maximal minimum distance of all known $q$-ary linear codes with relevant length and dimension, and its value is from [18].
@@
$q$ $p$ $b$ $r$ Lower bound Bose distance Parameter Optimality References
22 22 33 11 22 33 $17,8,6]^{*}$ optimal Theorem 5.1
22 22 33 11 22 33 $[17,9,5]$ optimal Theorem 5.1
22 33 22 11 88 $11$ $[73,37,13]$ $d_{\text{best}}=14$ Theorem 5.1
44 22 22 33 33 77 $17,8,8]^{*}$ best-known Theorem 5.1
44 22 22 33 22 77 $[17,9,7]$ best-known Theorem 5.1
55 22 22 22 44 55 $26,13,8]^{*}$ $d_{\text{best}}=10$ Theorem 5.2
55 22 22 44 44 99 $26,12,10]^{*}$ $d_{\text{best}}=11$ Theorem 5.3
55 22 22 44 33 44 $[26,14,8]$ $d_{\text{best}}=9$ Theorem 5.3
77 22 22 22 77 99 $[50,26,14]$ $d_{\text{best}}=15$ Theorem 5.3
88 22 22 11 $10$ $22$ $65,32,22]^{*}$ best-known Theorem 5.1
@@ 6 Constacyclic codes with $n=\Phi_{p_{1}p_{2}}(q)$
Recall that $q$ is a prime power, and $r\mid q-1$. We assume that $p_{1},p_{2}$ are two distinct primes with $p_{1}<p_{2}$. In the following, we construct constacyclic codes with length
@@
Moreover, we only consider the case when the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}|$, has exactly two elements, and then we use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
@@
Let $p_{1},p_{2}$ be two distinct primes with $p_{1}<p_{2}$, and $n=\Phi_{p_{1}p_{2}}(q)$. Assume $p_{2}\nmid(q^{p_{1}}-1)/r$. Then, an infinite family of $q$-ary $n,n-1-p_{1}p_{2}\lceil(n-1)/(2p_{1}p_{2})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p_{1}p_{2})\rceil}{q-1}\right\rceil+\Lambda_{p_{1}p_{2}}]$ constacyclic codes is constructed, where $\Lambda_{p_{1}p_{2}}$ has been defined in Theorem 3.4. Moreover, if either $r\leq 3$ or $2p_{1}p_{2}(q-1)\geq qr$, an infinite family of $q$-ary $n,n-p_{1}p_{2}\lfloor(n-1)/(2p_{1}p_{2})\rfloor,\geq\lfloor\frac{q(n-1)}{2p_{1}p_{2}(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}|$. Since $n=\Phi_{p_{1}p_{2}}(q)$ and $r\mid q-1$, by Lemma 2.1 we have $l_{i}\mid p_{1}p_{2}$ for any $i\in Z_{n,r}$.
@@
$\equiv 1+1+\cdots+1+1\equiv p_{2}\pmod{q^{p_{1}}-1}$.
@@
$\gcd\left(\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1},\frac{q^{p_{1}}-1}{r}\right)=\gcd\left(p_{2},\frac{q^{p_{1}}-1}{r}\right)=1,$
where the last equality follows from the assumption $p_{2}\nmid(q^{p_{1}}-1)/r$. So, noticing $\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1}=n\cdot\frac{q^{p_{2}}-1}{q-1}$, we obtain
$\gcd(n,(q^{p_{1}}-1)/r)=1.$ (6.3)
@@
Next, we want to compute $N_{p_{2}}$. Similarly as deducing (6.3), we have
$\gcd\left(n,\frac{q^{p_{2}}-1}{r}\right)\leq\gcd\left(\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1},\frac{q^{p_{2}}-1}{r}\right)=\gcd\left(p_{1},\frac{q^{p_{2}}-1}{r}\right)\leq p_{1}.$
@@
Hence, we have $\{l_{i}:\,i\in Z_{n,r}\}=\{1,p_{1}p_{2}\}.$ Then, the desired results follow directly from Theorems 3.1 and 3.4. ∎
We remark that in Theorem 6.1, since $N_{1}>0$, by Remark 3.2 we know that the constacyclic codes there are monomially equivalent to some cyclic codes over $\mathbb{F}_{q}$.
We also remark that the condition $p_{2}\nmid(q^{p_{1}}-1)/r$ in Theorem 6.1 holds when $p_{1}\nmid p_{2}-1$. Indeed, if $p_{1}\nmid p_{2}-1$, then $q^{p_{1}}\not\equiv 1\pmod{p_{2}}$, and so $p_{2}\nmid q^{p_{1}}-1$.
We notice that some codes in Theorem 6.1 are better than BCH constacyclic codes; see Example 6.2.
Example 6.2.
Let $q=7$, $p_{1}=2$, $p_{2}=3$, $r=2$, and $n=43$. Then, the $7$-cyclotomic cosets in $Z_{43,2}$ are: $C_{1}=\{1,7,37,49,79,85\}$, $C_{3}=\{3,21,25,61,65,83\}$, $C_{5}=\{5,13,35,51,73,81\}$, $C_{9}=\{9,11,23,63,75,77\}$, $C_{15}=\{15,19,39,47,67,71\}$, $C_{17}=\{17,27,33,53,59,69\}$, $C_{29}=\{29,31,41,45,55,57\}$, $C_{43}=\{43\}$. So, $\{l_{i},i\in Z_{43,2}\}=\{1,6\}$, $N_{1}=1$ and $N_{6}=7$. From the first family of constacyclic codes in Theorem 6.1, we take the defining set as the set $C_{1}\cup C_{3}\cup C_{5}\cup C_{9}\cup C_{43}$ and then obtain a $7$-ary constacyclic $[43,18,17]$ code, which is not a BCH constacyclic code ($d_{\text{best}}=18$ according to [18]). In this case, there is only one BCH constacyclic code of length 43 and dimension 18 (its defining set is $C_{1}\cup C_{5}\cup C_{15}\cup C_{29}\cup C_{43}$ and minimum distance is 16).
In Table LABEL:tab:p1p2, we present some codes constructed in Theorem 6.1 having good parameters.
@@
$q$ $p_{1}$ $p_{2}$ $r$ Lower bound Bose distance Parameter Optimality
@@
7 Concluding remarks
We constructed several infinite families of constacyclic codes over $\mathbb{F}_{q}$ in this paper. They contain many constacyclic codes with optimal, or almost-optimal, or best-known parameters. We presented some of them in Tables LABEL:tab:qp, LABEL:tab:qpb and LABEL:tab:p1p2, and the codes in these tables are new compared to the constacyclic codes in the references below. Especially, we considered two forms of the code length: $n=\Phi_{p^{b}}(q)$ or $\Phi_{p_{1}p_{2}}(q)$, which are new up to our knowledge. We also presented several codes which are better than BCH constacyclic codes (see Examples 4.2, 5.4 and 6.2).
For the construction, we first classified the cyclotomic cosets according to their sizes, and then we took the defining set as the union of the first half of cyclotomic cosets in each class (sorted in ascending order by their coset leaders). In this paper, we focused on the case when there are exactly two kinds of cyclotomic cosets. According to our Magma experimental data, when there are more than two kinds of cyclotomic cosets, there are still some good codes under the above construction.
We emphasize that the codes we constructed are not always BCH constacyclic codes. The other fact we want to emphasize is that if $N$ is the number of cyclotomic cosets with maximum size, then in the above construction, the first $\lfloor N/2\rfloor$ consecutive elements may be not always in the defining set (see Example 2.14). This might deserve further investigation.
For the specific codes presented in Sections 4, 5 and 6, except the codes in Theorem 5.3, the others have been proved to be monomially equivalent to some cyclic codes over $\mathbb{F}_{q}$. By looking at the defining sets of these cyclic codes, we haven’t found any simple rule which can give an infinite family of cyclic codes.
We also remark that for the second family of constacyclic codes in Theorems 4.1, 5.1, 5.2 and 6.1, if furthermore $r\leq 2$, then by Theorem 3.4, one can get a better lower bound on their minimum distances.